# Module-conditioned distribution of quantum circuits

### v2

@@ Abstract
As quantum computers require highly specialized and stable environments to operate, expanding their capabilities within a single system presents significant technical challenges. By interconnecting multiple quantum processors, distributed quantum computing can facilitate the execution of more complex and larger-scale quantum algorithms. End-to-end heuristics for the distribution of quantum circuits have been developed so far. In this work, we derive an exact integer programming approach for the Distributed Quantum Circuit (DQC) problem, assuming fixed module allocations. Since every DQC algorithm necessarily yields a module allocation function, our formulation can be integrated with it as a post-processing step. This improves on the hypergraph partitioning formulation, which finds a module allocation function and an efficient distribution at once. We also show that a suboptimal heuristic to find good allocations can outperform previous methods. In particular, for quantum Fourier transform circuits, we conjecture from experiments that the optimal module allocation is the trivial one found by this method.
@@
Distributed quantum computing, Integer programming
@@ 1 Introduction
In recent years, the capacity of quantum computing hardware has steadily increased [3, 6, 23, 19, 21]. However, a significant obstacle to realizing the full potential of quantum computing is the limited number of qubits available in a single quantum computer (or module) [29, 16, 25, 5]. To overcome thislimitation, distributing a large quantum computation over a network of modules has emerged as a promising approach [8, 7, 12, 33, 14].
@@
The Distributed Quantum Circuit (DQC) problem involves dividing a quantum circuit across multiple modules. A non-local controlled unitary gate between two modules can be realized from local operations if the modules share a Bell state [24, 28, 34], which is referred to as an ebit [17, 2, 32, 1]. Typically, DQC attempts to minimize the ebit cost for circuit execution. Its goal is to map the qubits in the circuit to individual modules and determine how to implement the non-local operations in a way that minimizes the number of ebits and communications.
@@
This paper primarily considers homogeneous networks with $k\geq 3$ modules. In this setting, distributions using a hypergraph partitioning formulation and heuristic solvers proved to be the most effective [32, 1]. However, this approach jointly determines the module allocation and the implementation of non-local gates based on that allocation, even though the former cannot logically depend on the latter. As our main contribution, we observe that exact integer programming formulations [26, 20, 31] yield improved ebit costs if the module allocation function is fixed, demonstrating stability for various types of circuits. First, we present formulations for $k\geq 4$ modules and then show how it simplifies for $k=3$ modules. As a result, our method can be combined with any DQC solver by ignoring its distribution steps and adopting only the module allocation it returns. Our approach further reduces ebit costs for various types of circuits, with significant improvements observed in some structured circuits. This highlights the importance of our post-processing for achieving robust results.
We also propose a simple suboptimal heuristic to find a module allocation function, although we currently find it effective only for quantum Fourier transform (QFT) circuits [24, 10, 27]. Interestingly, experimental results suggest that the proposed heuristic is likely to be optimal for these circuits.
@@
The rest of this paper is organized as follows. Section 2 provides an overview of background and related works. In Section 3, we present detailed derivations and proofs of our binary integer programming (BIP) formulations, along with an explanation of the unique characteristics of distributing QFT circuits. Section 4 presents experimental results, which have been validated on various circuits. Finally, Section 5 concludes the paper and offers a perspective on future work.
@@ 2.1 Distribution of quantum circuits
In the context of distributed quantum computing, a module refers to an individual quantum computer that forms part of a larger, interconnected network [1].
As in previous studies, we consider a quantum circuit to be composed only of unary and binary gates [17, 2, 32, 1]. Here, each binary gate is a controlled-phase ($CP$) gate [24]
@@
where $\mathbb{I}$ denotes the single-qubit identity operator. Any circuit can be transformed into an equivalent circuit in this form, a fact that follows from the universality of the set consisting of all unary gates and the CNOT gate [4].
Distributing an $n$-qubit quantum circuit involves two steps. First, each qubit must be assigned to a module. Second, non-local gates are identified based on this module allocation, and the optimal method to implement these gates with the least resources is determined. This overall process is referred to as the DQC problem. Typically, a $(k,\epsilon)$ balanced partition is used for module allocation, ensuring that each module contains at most $(1+\epsilon)\frac{n}{k}$ qubits [17]. By a reduction from the hypergraph min-cut problem, the DQC problem was shown to be NP-hard [2].
Figure 1: A non-local controlled unitary gate can be realized using only local operations and classical communications at the cost of a single ebit.
2.2 Cat-entanglement & cat-disentanglement
@@
$\frac{|00\rangle_{AB}+|11\rangle_{AB}}{\sqrt{2}}$
between two modules $A$ and $B$, which is a maximally entangled bipartite state. In the context of DQC, such states are referred to as ebits. Fig. 1 shows how to implement a non-local controlled unitary gate with a single control qubit $|c\rangle$ and a single target qubit $|t\rangle$ at the cost of a single ebit. The steps before (after) the controlled unitary gate constitute a cat-entanglement (cat-disentanglement) process. The former creates a linked copy of $|c\rangle$ in module $B$, and the latter effectively reverses this process. After cat-disentanglement, the operation $|ct\rangle\mapsto CP(\theta)|ct\rangle$ has been successfully implemented.
Suppose we create linked copies of a qubit $q$ at modules other than the one including $q$. This cat-entanglement operation (a) commutes with $CP$ gates acting on $q$ and (b) generally does not commute with arbitrary unary gates acting on $q$ (see appendices A & B). The first property allows us to restrict the creation of a linked copy of $q$ in any module to occur immediately after one of the unary operations on $q$ or at time 00, assuming that the circuit is compiled exclusively with $CP$ gates and unary gates. The second property implies that in general, all linked copies of $q$ in other modules must be reintegrated into $q$ before a unary gate acts on $q$.
@@ 2.3 Problem formulation
We represent a quantum circuit acting on a set of $n$ qubits $Q=\{q_{i}\}_{i=1}^{n}$ as a finite set $C=U\cup B$, where
@@
Figure 2: (a) Distribution of a quantum circuit and (b) its hypergraph representation. Gray boxes represent module indices and gray arrows indicate migrations.
is a set of non-local binary gates. As is clear from the notation, $(q_{i},t)$ represents a unary gate acting on $q_{i}$ at time $t$ and ($(\{q_{i},q_{j}\},t)$ represents a binary gate acting on $\{q_{i},q_{j}\}$ at time $t$. Without loss of generality, it is assumed that all $T^{(i)}$ and $T^{(\{i,j\})}$ are mutually exclusive.
In what follows, we assume that there are $k\geq 3$ modules in total, i.e., $P=\{p_{1},p_{2},\cdots,p_{k}\}$, and that we are given a module allocation function $\pi:Q\rightarrow P$. This implies that $\pi(q_{i})\neq\pi(q_{j})$ for all ($(\{q_{i},q_{j}\},t)\in B$. The concept of migration was first defined in Ref. [17], which essentially represents the cat-entanglement operation in Section 2.2. It implicitly assumes that cat-disentanglements are performed whenever necessary.
Definition 1 (Migration [17]).
@@
Definition 2 (Coverage [17]).
For a non-local $CP$ gate $g=(\{q_{i},q_{j}\},t^{*})$, we say that a migration $(q,p,t)$ (see 1 & 2) or a pair of migrations $\{(q_{i^{\prime}},p,t),(q_{j^{\prime}},p,t^{\prime})\}$ (see 3) covers $g$ if one of the following conditions holds, thereby enabling the execution of $g$.
@@
    $q=q_{i},\,p=\pi(q_{j}),\,t=\max\limits_{\tilde{t}\in\{0\}\cup T^{(i)},\,\tilde{t}\leq t^{*}}\tilde{t}$
@@
    $q=q_{j},\,p=\pi(q_{i}),\,t=\max\limits_{\tilde{t}\in\{0\}\cup T^{(j)},\,\tilde{t}\leq t^{*}}\tilde{t}$
  3. 3.
    $\{i^{\prime},j^{\prime}\}=\{i,j\},\,p\notin\{\pi(q_{i}),\pi(q_{j})\},\\ t=\max\limits_{\tilde{t}\in\{0\}\cup T^{(i^{\prime})},\,\tilde{t}\leq t^{*}}\tilde{t},\,t^{\prime}=\max\limits_{\tilde{t}\in\{0\}\cup T^{(j^{\prime})},\,\tilde{t}\leq t^{*}}\tilde{t}$
Given $\pi$, the problem of finding the minimum subset of $M$ that covers all gates in $B$ following Definition 2 is called migration selection under general coverage (MS-GC). If we do not allow 3 in Definition 2, then the problem is called migration selection under home coverage (MS-HC). An efficient optimal algorithm for MS-HC given $\pi$ has been identified. On the other hand, MS-GC is conjectured to be NP-hard, and a heuristic was proposed, which greedily adds a set of migrations at each iteration based on some criterion [17].
@@ 2.4 Hypergraph partitioning formulation
The problem of quantum circuit distribution can be reduced to a hypergraph partitioning problem [2], for which efficient heuristic solvers exist. This currently achieves the best results for $k\geq 3$ [32, 1]. Given a circuit $C$, the corresponding hypergraph is constructed as follows. First, a node is added for each binary gate and for each qubit. Second, for each qubit $q_{i}$ and time interval $I=(t,t^{\prime})$ such that the following conditions are satisfied, a hyperedge that connects $q_{i}$ and all the binary gates that act on $q_{i}$ within $I$ is added, if such gate exists:
@@
Fig. 2 shows an example circuit distributed by this method. The coverage of the gate $\alpha$ corresponds to 3 in Definition 2, and the coverages of the gates $\beta$ and $\gamma$ corresponds to 1 or 2 in Definition 2.
3 Main results

### v3

# Binary integer programming for optimizing ebit cost in distributed quantum circuits with fixed module allocation

@@ Abstract
Modular and networked quantum architectures can scale beyond the qubit count of a single device, but executing a circuit across modules requires implementing non-local two-qubit gates using shared entanglement (ebits) and classical communication, making ebit cost a central resource in distributed execution. The resulting distributed quantum circuit (DQC) problem is combinatorial, motivating prior heuristic approaches such as hypergraph partitioning. In this work, we decouple module allocation from distribution. For a fixed module allocation (i.e., assignment of each qubit to a specific Quantum Processing Unit), we formulate the remaining distribution layer as an exact binary integer programming (BIP). This yields solver-optimal distributions for the fixed-allocation subproblem and can be used as a post-processing step on top of any existing allocation method. We derive compact BIP formulations for four or more modules and a tighter specialization for three modules. Across a diverse benchmark suite, BIP post-processing reduces ebit cost by up to 20% for random circuits and by more than an order of magnitude for some arithmetic circuits. While the method incurs offline classical overhead, it is amortized when circuits are executed repeatedly.
@@
Distributed quantum computing, Cat-entanglement, Integer programming, Hypergraph partitioning
@@ 1 Introduction
In recent years, the capacity of quantum computing hardware has steadily increased [arute2019quantum, bravyi2018quantum, neill2018blueprint, kjaergaard2020superconducting]. However, a significant obstacle to realizing the full potential of quantum computing is the limited number of qubits available in a single quantum computer (or module) [terhal2015quantum, fowler2012surface, o2016scalable, bourassa2021blueprint]. To overcome thislimitation, distributing a large quantum computation over a network of modules has emerged as a promising approach [cirac1999distributed, cacciapuoti2019quantum, cuomo2023optimized, yimsiriwattana2004generalized, eisert2000optimal].
@@
The distributed quantum circuit (DQC) problem involves dividing a quantum circuit across multiple modules. A non-local controlled unitary gate between two modules can be realized using only local quantum operations and classical communication if the modules share a Bell state [nielsen2010quantum, sych2009complete, zaman2018counterfactual], which is referred to as an ebit [g2021efficient, andres2019automated, wu2023entanglement, andres2024distributing]. Typically, an algorithm used to solve DQC attempts to minimize the ebit cost for circuit execution. Its goal is to map the qubits in the circuit to individual modules and determine how to implement the non-local operations between modules in a way that minimizes the number of required ebits.
@@
This paper primarily considers homogeneous networks with $k\geq 3$ modules. In this setting, distributions using a hypergraph partitioning formulation and heuristic solvers proved to be the most effective [wu2023entanglement, andres2024distributing]. However, this approach jointly determines the module allocation and the implementation of non-local gates based on that allocation, even though the former cannot logically depend on the latter. As our main contribution, we observe that exact binary integer programming (BIP) formulations [papadimitriou1998combinatorial, karp2009reducibility, wolsey2020integer] under a fixed module allocation function (i.e., assignment of each qubit to a specific Quantum Processing Unit (QPU)) yield improved ebit costs, demonstrating stability for various types of circuits. First, we present formulations for $k\geq 4$ modules and then show how it simplifies for $k=3$ modules. As a result, our method can be combined with any DQC solver by ignoring its distribution steps and adopting only the module allocation it returns. Our approach further reduces ebit costs for various types of circuits, with significant improvements observed in some structured circuits. This highlights the importance of our post-processing for achieving robust results.
The motivation for using a BIP formulation is that, given a fixed module allocation, selecting migrations (defined in Section 2.3) to cover all non-local gates reduces to a discrete subset-selection problem with binary decisions. An exact BIP formulation yields (solver-)optimal solutions for this fixed-module subproblem, enabling a principled post-processing step that can be plugged into any existing method by reusing its module allocation and optimizing only the distribution layer.
@@
The rest of this paper is organized as follows. Section 2 provides an overview of the background and related works. In Section 3, we present detailed derivations and proofs of our BIP formulations, along with an explanation of the unique characteristics of distributing quantum Fourier transform (QFT) circuits [nielsen2010quantum, coppersmith2002approximate, ruiz2017quantum]. Section 4 presents experimental results, which have been validated on various circuits. Finally, Section 5 concludes the paper and offers a perspective on future work.
@@ 2 Preliminaries
At its core, the objective of this work can be summarized as follows: Given a monolithic quantum circuit $C$, how can we optimally distribute its constituent gates across $k$ distinct hardware modules while minimizing the overhead of entanglement resources and non-local operations? In the following sections, we transition into the formal definitions of distributed quantum architectures. The concepts introduced in this section are used throughout and will be required to state our BIP formulation in Section 3.
Figure 1: Example network of five QPUs. Every pair of modules shares entanglement resources to facilitate non-local gate operations. This distributed architecture enables large-scale quantum circuits through a combination of local quantum operations and inter-module classical communication.
@@ 2.1 Distribution of quantum circuits
In the context of distributed quantum computing, a module refers to an individual quantum computer that forms part of a larger, interconnected network [andres2024distributing] (represented as a box in Fig. 1). Modern quantum architectures often face scaling limitations where the qubit count of a single device is insufficient for large-scale algorithms. To overcome this, software tools must partition a global circuit into smaller fragments that can be executed across multiple modules, allowing them to function collectively as a single, powerful processor. This distributed execution, however, necessitates sophisticated non-local gate operations and advanced compiler strategies to determine an optimal circuit layout, which amounts to minimizing the entanglement and communication overhead between modules.
As in previous studies, we consider a quantum circuit to be composed only of unary and binary gates [g2021efficient, andres2019automated, wu2023entanglement, andres2024distributing]. Here, each binary gate is a controlled phase ($CP$) gate [nielsen2010quantum]
@@
where $\mathbb{I}$ denotes the single-qubit identity operator. Any circuit can be transformed into an equivalent circuit in this form, a fact that follows from the universality of the set consisting of all unary gates and the CNOT gate [barenco1995elementary].
Distributing an $n$-qubit quantum circuit involves two steps. First, each qubit must be assigned to a module. Second, non-local gates are identified based on this module allocation, and the optimal method to implement these gates with the least resources is determined. This overall process is referred to as the DQC problem. Typically, a $(k,\epsilon)$ balanced partition is used for module allocation, ensuring that each module contains at most $(1+\epsilon)\frac{n}{k}$ qubits [g2021efficient]. By a reduction from the hypergraph min-cut problem, the DQC problem was shown to be NP-hard [andres2019automated]. This complexity implies that finding a globally optimal distribution for large-scale circuits is computationally intractable, traditionally necessitating the use of heuristic approaches such as hypergraph partitioning. However, these heuristics often bundle module allocation and ebit distribution together, potentially missing significant resource savings. By decoupling these steps, we can isolate the ebit distribution layer as a subproblem that can be solved via BIP. This allows our method to refine heuristic-based pipelines, providing a better solution for a given module allocation under which a standard heuristic alone would settle for suboptimal resource costs.
Figure 2: Implementation of a non-local gate. Module AA contains qubits $(a,e)$ and Module $B$ contains qubits $(e^{\prime},b)$. The controlled-$U$ gate $CU_{a\to b}$ with control $a$ and target $b$ can be realized using only local operations and classical communication at the cost of a single ebit $|\Phi^{+}\rangle_{ee^{\prime}}$ shared between AA and $B$.
2.2 Cat-entanglement and cat-disentanglement
@@
$|\Phi^{+}\rangle_{AB}=\frac{|00\rangle_{AB}+|11\rangle_{AB}}{\sqrt{2}}$
between two modules $A$ and $B$, which is a maximally entangled bipartite state. In the context of DQC, such states are referred to as ebits. Fig. 2 shows how to implement a non-local controlled unitary gate with a single control qubit $a$ and a single target qubit $b$ at the cost of a single ebit (for a complete validation of this protocol, see Appendix A). The steps before (after) the controlled unitary gate constitute a cat-entanglement (cat-disentanglement) process. The former creates a linked copy of $a$ in module $B$, and the latter effectively reverses this process. After cat-disentanglement, operations such as $CP(\theta)_{a\to b}$ can be successfully implemented. In practice, the end-to-end fidelity and runtime of this ebit-assisted cat-(dis)entanglement are currently limited by the mid-circuit measurements and real-time feed-forward, but existing demonstrations already mitigate part of this overhead to make non-local operations more robust [carrera2024combining, baumer2025measurement].
Suppose we create linked copies of a qubit $q$ at modules other than the one including $q$. This cat-entanglement operation (i) commutes with $CP$ gates acting on $q$ and (ii) generally does not commute with arbitrary unary gates acting on $q$ (see Appendices B and C). The first property allows us to restrict the creation of a linked copy of $q$ in any module to occur immediately after one of the unary operations on $q$ or at time 0, assuming that the circuit is compiled exclusively with $CP$ gates and unary gates. The second property implies that in general, all linked copies of $q$ in other modules must be reintegrated into $q$ before a unary gate acts on $q$.
@@ 2.3 Problem formulation
To formally address the distribution of gates across a modular architecture, we first define the structure of the input quantum circuit $C$. Our goal is to transform this monolithic circuit into a partitioned representation that can be mapped onto $k$ hardware modules. To achieve this, we model the circuit as a collection of discrete gate events in space and time, categorized by their qubit dependencies.
Specifically, we represent a quantum circuit acting on a set of $n$ qubits $Q=\{q_{i}\}_{i=1}^{n}$ as a finite set $C=U\cup B$, where
@@
is a set of non-local binary gates. As is clear from the notation, $(q_{i},t)$ represents a unary gate acting on $q_{i}$ at time $t$ and $(\{q_{i},q_{j}\},t)$ represents a binary gate acting on $\{q_{i},q_{j}\}$ at time $t$. Without loss of generality, it is assumed that all $T^{(i)}$ and $T^{(\{i,j\})}$ are mutually exclusive.
In what follows, we assume that there are $k\geq 3$ modules in total, i.e., $P=\{p_{1},p_{2},\cdots,p_{k}\}$, and that we are given a module allocation function $\pi:Q\rightarrow P$. This implies that $\pi(q_{i})\neq\pi(q_{j})$ for all $(\{q_{i},q_{j}\},t)\in B$. The concept of migration was first defined in Ref. [g2021efficient]. Intuitively, a migration does not physically move a qubit between modules. Rather, it denotes a temporary, protocol-level operation in which the quantum information of a logical qubit is made available at another module so that subsequent $CP$ gates that would otherwise be non-local can be executed locally at that module. Technically, a migration represents the cat-entanglement operation in Section 2.2. It also implicitly assumes that cat-disentanglements are performed whenever necessary.
Definition 1 (Migration [g2021efficient]).
@@
For a non-local $CP$ gate, home coverage restricts us to migrating one endpoint qubit to the home module of the other endpoint, so that the gate can be executed where one of its qubits already belongs. In contrast, joint coverage allows two migrations executing the gate at a third module by migrating both endpoint qubits there. This greatly enlarges the combinatorial choices because it couples the selection decisions for the two qubits. Formally, we have the following definitions.
Definition 2 (Home coverage [g2021efficient]).
For a non-local $CP$ gate $g=(\{q_{i},q_{j}\},t^{*})$, we say that a migration $(q,p,t)$ home-covers $g$ if one of the following conditions holds, thereby enabling the execution of $g$.
@@
    $q=q_{i},\quad p=\pi(q_{j}),\quad t=\max\limits_{\tilde{t}\in\{0\}\cup T^{(i)},\,\tilde{t}\leq t^{*}}\tilde{t}$
@@
    $q=q_{j},\quad p=\pi(q_{i}),\quad t=\max\limits_{\tilde{t}\in\{0\}\cup T^{(j)},\,\tilde{t}\leq t^{*}}\tilde{t}$
Definition 3 (Joint coverage [g2021efficient]).
For a non-local $CP$ gate $g=(\{q_{i},q_{j}\},t^{*})$, we say that a pair of migrations $\{(q_{i^{\prime}},p,t),(q_{j^{\prime}},p,t^{\prime})\}$ joint-covers $g$ if the following condition holds, thereby enabling the execution of $g$.
  • $\{i^{\prime},j^{\prime}\}=\{i,j\},\quad p\notin\{\pi(q_{i}),\pi(q_{j})\},\\ t=\max\limits_{\tilde{t}\in\{0\}\cup T^{(i^{\prime})},\,\tilde{t}\leq t^{*}}\tilde{t},\qquad t^{\prime}=\max\limits_{\tilde{t}\in\{0\}\cup T^{(j^{\prime})},\,\tilde{t}\leq t^{*}}\tilde{t}$
Given $\pi$, the problem of finding the minimum subset of $M$ that covers all gates in $B$ following Definition 2 is called migration selection under home coverage (MS-HC). If we further allow joint coverage in Definition 3, then the problem is called migration selection under general coverage (MS-GC). We remark that MS-HC can be viewed as a set cover instance with highly restricted structure, and an optimal polynomial-time algorithm for MS-HC given $\pi$ has been identified. On the other hand, joint coverage significantly expands the search space, and MS-GC is conjectured to be NP-hard. A heuristic for MS-GC was proposed, which greedily adds a set of migrations at each iteration based on some criterion [g2021efficient].
Importantly, this work does not claim a polynomial-time algorithm for MS-GC. Instead, our contribution is an exact BIP formulation for MS-GC under a fixed $\pi$, which we use as a post-processing step. Starting from any candidate allocation $\pi$ (e.g., produced by a heuristic allocator), we solve the corresponding fixed-$\pi$ BIP to obtain an ebit distribution for that allocation. This improves the ebit cost without changing the module allocation.
@@ 2.4 Hypergraph partitioning formulation
Figure 3: (a) Example distributed circuit for $k=3$ modules. Wires with the same color belong to the same module. The labels $1,1,2,2,3,3$ on the left are the module indices of the six logical qubits (each module contains two qubits). White squares represent single-qubit gates and the symbols $\alpha,\beta,\gamma,\delta$ denote two-qubit gates (with repeated $\delta$-blocks). Downward arrows indicate migrations (here, to module 3). (b) Hypergraph representation of the same instance, with qubit vertices $q_{1},\ldots,q_{6}$ and gate vertices $\alpha,\beta,\gamma,\delta$. Qubit-vertex colors indicate their home modules, whereas gate-vertex colors indicate the modules where the corresponding gates are executed.
Our BIP post-processing requires a module allocation $\pi$ as input. In practice, we can obtain $\pi$ from existing allocation heuristics, and in particular from the hypergraph partitioning formulation of circuit distribution [andres2019automated], which has achieved the best performance for $k\geq 3$ [wu2023entanglement, andres2024distributing]. We briefly summarize this construction here for completeness. Readers interested only in the post-processing formulation may skip to Section 3.
Given a circuit $C$, the corresponding hypergraph is constructed as follows. First, a node is added for each binary gate and for each qubit. Second, for each qubit $q_{i}$ and time interval $I=(t,t^{\prime})$ such that the following conditions are satisfied, a hyperedge that connects $q_{i}$ and all the binary gates that act on $q_{i}$ within $I$ is added, if such gate exists:
@@
Fig. 3 shows an example circuit distributed by this method. The coverage of the gate $\alpha$ is a joint coverage, and the coverages of the gates $\beta$ and $\gamma$ are home coverages.
3 Algorithmic improvements and examples
The full DQC pipeline consists of two layers: a module allocation step that fixes $\pi:Q\to P$, followed by a distribution step that selects migrations to cover all non-local gates. The latter step is precisely MS-GC under a fixed $\pi$, and existing approaches rely on heuristics because MS-GC is conjectured to be NP-hard. Our goal in this section is therefore not to obtain a new allocator, but to provide a post-processing method that, for any given $\pi$, computes the set of migrations for that allocation. To this end, we formulate fixed-$\pi$ MS-GC as a BIP, which serves as a drop-in improvement layer on top of any module allocation heuristic by optimizing only the migration-selection layer. In Section 3.3, we include a small illustrative example showing that even on a tiny circuit the hypergraph partitioner can be suboptimal, and that our BIP post-processing reduces the ebit cost.
We begin by describing a BIP formulation of MS-GC for $k\geq 4$. A more efficient formulation for $k=3$ is described in Section 3.2.